Beau Glaser wishes to strengthen a mixture from 10% alcohol to 30% alcohol. How much pure alcohol should be added to 7 L of the 10% mixture?
Answer by checkley77(12844)

.10*7+x=.30(7+x)
.7+x=2.1+.3x
x-.3x=2.1-.7
.7x=1.4
x=1.4/.7
x=2 liters of pure alcohol is needed.
Proof:
.10*7+2=.30(7+2)
.7+2=.3*9
2.7=2.7